what is the equation of x/2+3/8=1

Answer by jndarrell(58)

You need to find the LCD (lowest common denominator). In this case it is 8. So you will take the 8 and multiply all terms on both side of the equation by 8.
Don't forget the 1. what you do to one side, you must do to both.
8(x/2) + 8(3/8) = 1(8)
8(x/2)..the two cancels out and you are left with 4x
8(3/8)..the 8 cancels and leaves with 3
1(8)=8
4x+3=8
Subtract the 3 from both sides to get the x term alone
4x=5
Divide both sides by x to get alone
x=5/4
:)